OpenStudy (anonymous):

True or False: The following function is invertible: f(x) = x2 Select one: True False

OpenStudy (anonymous):

i think false

OpenStudy (anonymous):

f(x)=x^2 y=x^2 x=y^2 sqrtx=y

OpenStudy (help!!!!):

its \[x ^{2}\] right?

OpenStudy (anonymous):

yes

OpenStudy (help!!!!):

Use the horizontal test, if there are more than 2 point than it is uninvertible

OpenStudy (anonymous):

so it does have an inverce

OpenStudy (help!!!!):

wait a sec it doesn't. x^2 does hit more than one point. I made a type earlier from 2 point to 1 point